OpenStudy (barrelracing):

Which of the following expressions represents a function? a) x=1 b) {(3,2),(3,-2),(4,5),(4,-5) c) y=4x-1 d) 4x^2 + y^2 = 16

OpenStudy (anonymous):

function meaning graph?

OpenStudy (barrelracing):

not exactly sure

OpenStudy (anonymous):

a c d are graphs.b is coordinates

OpenStudy (anonymous):

By the way you look like Kim Kardashian in the photo

OpenStudy (haseeb96):

what do you think which one is correct ?

OpenStudy (barrelracing):

@jim_thompson5910 im not sure about any of this

OpenStudy (anonymous):

C, because a function by definition is one where the there is only 1 solution (y value) for an x coordinate. x=1 is a 1 to many, d is many to many, and b is a list i think.

OpenStudy (haseeb96):

f(x) = x^2 +1 where y=f(x) this type of equation is called function now guess what is the correct answer

OpenStudy (haseeb96):

@Barrelracing tell me ?

OpenStudy (barrelracing):

c

OpenStudy (haseeb96):

yes correct

jimthompson5910 (jim_thompson5910):

Let's go through the choices a) x=1 this is NOT a function because it plots a vertical line. All vertical lines are NOT functions. ------------------------------------------------------- b) {(3,2),(3,-2),(4,5),(4,-5) This is NOT a function either. The points (3,2) and (3,-2) violate the vertical line test. ------------------------------------------------------- c) y=4x-1 This is a function because each x value plugged in produces exactly one y value. The graph passes the vertical line test because any given vertical line only passes through exactly 1 point. ------------------------------------------------------- d) 4x^2 + y^2 = 16 This makes a circle. It is possible to draw a vertical line through more than one point on this circle. So it fails the vertical line test and it is NOT a function.

OpenStudy (anonymous):

Well technically d is an elipse...

jimthompson5910 (jim_thompson5910):

oh right, my bad